what is the strength of an electric field that will balance the weight of a 3.5 g plastic sphere that has been charged to -8.4 nC?
Answer #1

Gravitational Force = Electrostatic Force
m*g = q*E
3.5 * 10^-3 * 9.8 = 8.4 * 10^-9 * E
E =  4.1 * 10^6 N/C

Strength of an electric field , E =  4.1 * 10^6 N/C